# Attention-Guided Multi-scale Interaction Network for Face Super-Resolution

### v3

@@ III-B3 Selective Kernel Attention Fusion Module (SKAF)
Inspired by SKNet [36] and LSKNet [37], as shown in Fig. 2, we design an SKAF module to give our model the ability to select local and global features required for reconstruction for fusion interaction. Specifically, SKAF first extracts both local and global features using convolutions with different receptive fields. It then computes feature importance through an average pooling operation, which was previously not explicitly shown in the diagram. Finally, an adaptive weighting mechanism selects and fuses the most relevant local and global information, enabling SKAF to dynamically emphasize critical features under varying receptive fields. This design enhances the model’s capacity for feature selection, improving both its performance and generalization in face super-resolution. Given feature obtained by the SA and the RDFE, we first fuse the local and global features extracted by a convolution and a convolution to get a hybrid feature . This operation can be expressed as
@@
where , and indicate the average and max pooling operation along the channel direction and sigmoid function, respectively. Finally, we multiply the weights obtained from the above calculations with the local and global features, respectively. Thus, our SKAF can adaptively select the important local and global information required for reconstruction. The process of obtaining important local and global features , by adaptive weight selection can be expressed as
@@ III-C Encoder and Decoder Feature Fusion Module (EDFF)
To fully utilize the multi-scale features extracted from the encoding and decoding stage, we introduce an EDFF to fuse different features, enabling our AMINet with better feature propagation and representation capabilities. As shown in Fig. 2, our EDFF mainly utilizes our proposed SKAF to fuse and select different scale features required for reconstruction. Given the feature , the feature from the decoding stage and the encoding stage, respectively. Firstly, we concatenate features from the encoding and decoding stages along the channel dimension. Then, a convolution is used to reduce the channel counts and reduce the process’s computational costs to obtain two weights through our SKAF, which can be expressed as
where represents Selective Kernel Attention Fusion Module, stands for the convolutional layer, and denotes the operation of concatenating features across the channel dimension. Next, we feed the two obtained weights into two branches for multiplication. Through this operator, we obtain the selected facial features from hybrid features obtained by fusing the encoding and decoding features. Finally, we add the features of the two branches. The process is denoted as
@@
| Methods | PSNR | SSIM | VIF | LPIPS |
| LGFI w/o SA | 27.75 | 0.7944 | 0.4652 | 0.1886 |
| LGFI w/o RDFE | 27.51 | 0.7840 | 0.4495 | 0.2085 |
| LGFI w/o SKAF | 27.74 | 0.7932 | 0.4611 | 0.1979 |
| LGFI | 27.83 | 0.7961 | 0.4725 | 0.1821 |
@@ III-D2 Perceptual loss
To enhance the visual quality of super-resolution images, we apply perceptual loss. This involves using a pre-trained VGG19 [40] model to extract facial features from both the HR images and our generated FSR images. Then, we compare the obtained perceptual features of HR and FSR images to constrain the generation of FSR features. Therefore, the perceptual loss can be described as
@@ III-D3 Adversarial loss
GANs have been shown to be effective in reconstructing photorealistic images [38, 39]. GAN generates FSR results through the generator while using the discriminator to distinguish between ground truth and FSR results, which ultimately enables the generator to generate realistic FSR results in the process of constant confrontation. This process is denoted as
@@ IV Experiments
This section will introduce the experimental setting, our ablation studies, and comparisons with current advanced FSR methods to concretely demonstrate the advantages of our approach on both synthetic and real test datasets.
@@ IV-A Datasets and Evaluation Metrics
In our studies, we utilize the CelebA [4] dataset for training and evaluation on CelebA [4], Helen [41], and SCface [42] datasets, respectively. We center-crop the aligned face images and resize them to pixels to obtain high-resolution (HR) versions. These HR images are then downsampled to pixels using bicubic interpolation, producing the corresponding low-resolution (LR) images. For our experiments, we randomly chose 18,000 CelebA images for training and 1,000 for testing. In addition, we also utilize the SCface test set as a real-world evaluation dataset. To measure the quality of the FSR results, we use five metrics: PSNR [43], SSIM [43], LPIPS [44], VIF [45], and FID [46].
@@ IV-B Implementation details
We implement our model using the PyTorch framework on an NVIDIA GeForce RTX 3090. The network is optimized using the Adam optimizer, with parameters set to and . The initial learning rate is , with separate learning rates for the generator and discriminator set at and , respectively. The loss function weights are configured as , , and .
@@ IV-C1 Study of LGFI
LGFI is proposed to extract local features and global relationships of images, which represents a new attempt to interact with local and global information. To verify the reasonableness of our design of LGFI, as shown in Table I, we design four ablation models. The first model removes the SA, labeled “LGFI w/o SA”. The second model removes RDFE, labeled as “LGFI w/o RDFE”. The third model removes SKAF, labeled as “LGFI w/o SKAF”. We have the following observations: (a) Introducing SA and RDFE alone can improve model performance. This is because the above two modules can capture local and global features to promote facial feature reconstruction, including facial details and overall contours; (b) Model performance has been significantly increased by introducing the SKAF to capture the relationship between local and global facial features. This is because our SKAF can promote interaction between our SA and RDFE, integrating richer information and providing supplementary information for the final FSR.
Furthermore, we provide a visual comparison in Fig. 5, illustrating the impact of removing certain components from LGFI. The reconstructed images exhibit noticeable blurring or artifacts, highlighting the importance of these components. Additionally, Fig. 6 presents heatmap visualizations of the outputs from different components within LGFI, with their corresponding locations in the network indicated in Fig. 2. Specifically, SKAF effectively integrates the global contours captured by SA with the regional features extracted by RDEF. This integration enables the model to focus more on essential facial structures and components while reducing emphasis on less critical details, such as hair.
@@ IV-C2 Comparison between LGFI and Transformer
As shown in Fig. 4, LGFI uses a dual-branch structure to interact with representing the local and global features. In contrast, the traditional Transformer in Restormer [55] uses a serial structure to link the local and global features. To verify the effectiveness of LGFI, we replace all LGFIs in the network with Transformers and conduct comparative experiments with similar parameters between the two models. From Table II, we can see that the network’s performance using LGFI is better when the two networks maintain similar parameters. This is because LGFI utilizes the features of both local and global branches for interaction, facilitating the communication of multi-scale facial information.
@@ IV-C3 Comparison between RDFE and FFN
The feed-forward network (FFN) performs independent nonlinear transformations of the inputs at each position to help the Transformer capture local features, but it lacks the ability to extract multi-scale features, which is not favorable for accurate FSR. In contrast, our RDFE can extract multi-scale local features well. To compare RDFE and FFN, we replace RDFE with FFN while keeping the parameters of the two models similar. As shown in Table III, since FFN’s ability to capture feature interactions is limited compared to our RDFE which utilizes multiple branches to capture different receptive field facial features, our RDFE performs much better than FFN with similar computational consumption.
@@ IV-C4 Effectiveness of RDFE
In RDFE, a three-branch network guided by an attention mechanism is used for deep feature extraction, and the feature refinement module is used to enrich feature representation. To verify the effectiveness of RDFE, we conduct multiple ablation experiments. We designed five improved models. The first model adopts a single branch structure of depthwise convolution, labeled as “Single path ( dw)”. The second model adopts a single branch structure of depthwise convolution, labeled as “Single path ( dw)”. The third model adopts a single branch structure of depthwise convolution, labeled as “Single path ( dw)”. The fourth model removes attention units labeled as ”w/o AU”. The fifth model removes the feature refinement module, labeled as ”w/o FRM”. From the Table IV, we have the following observations: (a) By comparing the first three rows and the last row of the table, it can be seen that multi-scale branching facilitates the model’s performance due to its ability to extract face features at different levels; (b) From the comparison between the second and the last rows of the table and the last row, it can be seen that using attention units (AU) to guide three-branch feature extraction can enable the model to adaptively allocate weights, enhance the representation of important facial information, and thus improve model performance; (c) From the last two rows of the table, we can conclude that the feature refinement module (FRM) can further integrate multi-scale information, refine multi-scale fusion features, and thus improve performance.
@@ IV-C6 Study of EDFF
This section presents a set of experiments to validate the effectiveness of our EDFF, a module tailored for fusing multi-scale features. We add EDFF to SPARNet [17], which uses EDFF to connect the encoding and decoding stages in SPARNet and send them to the next decoding stage. Additionally, we add EDFF to SFMNet [20], and the specific operation is the same as in SPARNet. From the results of Table VI, we can see that although the parameters of both models increase slightly, the performance of the models improves, which precisely proves that EDFF is helpful for feature fusion in the encoding and decoding stages.
@@
| Methods | PSNR | SSIM | VIF | FID |
| FSRGAN [12] | 25.02 | 0.7279 | 0.3400 | 146.55 |
| DICGAN [14] | 25.59 | 0.7398 | 0.3925 | 144.25 |
| SPARGAN [17] | 25.86 | 0.7518 | 0.3932 | 149.54 |
| SFMGAN [20] | 25.96 | 0.7618 | 0.4019 | 141.23 |
| AMIGAN (Ours) | 26.35 | 0.7769 | 0.4101 | 122.43 |
@@ IV-D Comparisons with Other Methods
This section compares our AMINet and its GAN-based variant with leading FSR methods currently available, including SAN [47], RCAN [24], HAN [48], SwinIR [9], FSRNet [12], DICNet [14], FACN [49], SPARNet [17], SISN [18], AD-GNN [50], Restormer-M [51], LAAT [52], ELSFace [53], SFMNet [20] and SPADNet [54].
IV-D1 Comparison on CelebA dataset
@@
IV-D2 Comparison on Helen dataset
@@
IV-D3 Comparison with GAN-based methods
We present AMIGAN as an innovative approach to bolster the visual fidelity of image restoration tasks. To substantiate its superiority, we have conducted a rigorous comparison of AMIGAN against state-of-the-art GAN-based methodologies, namely FSRGAN [12], DICGAN [14], SPARGAN [17], and SFMGAN [20]. As a complementary assessment metric, we introduce the FID [46] to quantitatively evaluate the GANs’ performance. The outcomes presented in Table IX, derived from tests on the Helen dataset, reveal that AMIGAN outpaces its competitors considerably. Furthermore, the visual inspection illustrated in Fig. 10 underscores AMIGAN’s exceptional capabilities. Unlike existing FSR methods, which exhibit visible artifacts in generated facial images, AMIGAN meticulously restores critical facial features and intricate texture details around the mouth and nose, which underscores AMIGAN’s prowess in facial texture restoration, resulting in a notable enhancement in clarity and overall visual realism.
IV-D4 Comparison on Real-world surveillance faces
@@
All the above comparisons are tested on synthetic test sets, which fail to simulate real-world scenarios accurately. To further evaluate our model’s performance in real-world conditions, we also conduct experiments using low-quality face images from the SCface dataset [42]. As shown in Fig. 11, we compare the reconstruction results. From this figure, we find that the reconstruction results of face prior-based methods are not satisfactory. The challenge lies in accurately estimating priors from real-world LR facial images. Incorrect prior information can lead to misleading guidance during the reconstruction process. In contrast, our AMINet can restore clearer face details and faithful face structures. As shown in Table VIII, we also provide a quantitative comparison of cosine similarity using the above methods. This result fully demonstrates our method’s effectiveness in real scenarios.
IV-E Model Complexity Analysis
@@
In addition to the performance indicators mentioned earlier, the number of model parameters, inference time, and computational complexity are crucial factors in evaluating performance. As shown in the Table X, we have selected some models for comparison. Meanwhile, as shown in Fig. 1, we compare our model with existing ones in terms of parameters, PSNR values, and inference speed. We can see that AMINet can have faster inference time, smaller parameter count, and computational complexity with similar performance to SCTANet. In addition, we provide a small parameter version called ”AMINet-S”. AMINet-S performs similarly to SFMNet in terms of parameter quantity, while AMINet has more advantages in computational complexity and inference time. This is thanks to AMINet achieving high efficiency by promoting extensive multi-scale feature exchange within the network. This design enables features from different receptive fields to interact effectively, allowing the model to select the most relevant information for facial reconstruction adaptively. Therefore, AMINet maintains strong FSR performance while balancing parameters and computational costs.
@@ V Discussion and Future Works
Future work will focus on enhancing AMINet’s robustness to extreme poses and challenging lighting conditions while improving real-time inference efficiency. This includes integrating pose-invariant feature learning through attention-based mechanisms or 3D-aware priors for better side-profile reconstruction and developing adaptive illumination correction using physics-based relighting models or low-light enhancement techniques. Additionally, optimizing the model with lightweight architectures, quantization, and efficient inference strategies will enable faster inference while maintaining high-quality face reconstruction.

### v4

@@ III-B3 Selective Kernel Attention Fusion Module (SKAF)
Inspired by SKNet [35] and LSKNet [36], as shown in Fig. 2, we design an SKAF module to give our model the ability to select local and global features required for reconstruction for fusion interaction. Specifically, SKAF first extracts both local and global features using convolutions with different receptive fields. It then computes feature importance through an average pooling operation, which was previously not explicitly shown in the diagram. Finally, an adaptive weighting mechanism selects and fuses the most relevant local and global information, enabling SKAF to dynamically emphasize critical features under varying receptive fields. This design enhances the model’s capacity for feature selection, improving both its performance and generalization in face super-resolution. Given feature obtained by the SA and the RDFE, we first fuse the local and global features extracted by a convolution and a convolution to get a hybrid feature . This operation can be expressed as:
@@
where , , and indicate the average and max pooling operation along the channel direction and sigmoid function, respectively. Finally, we multiply the weights obtained from the above calculations with the local and global features, respectively. Thus, our SKAF can adaptively select the important local and global information required for reconstruction. The process of obtaining local and global features , by adaptive weight selection is:
@@ III-C Encoder and Decoder Feature Fusion Module (EDFF)
To fully utilize the multiscale features extracted from the encoding and decoding stage, we introduce an EDFF to fuse different features, enabling our AMINet with better feature propagation and representation capabilities. As shown in Fig. 2, our EDFF mainly utilizes our proposed SKAF to fuse and select different-scale features required for reconstruction. Given the feature , the feature is from the decoding stage and the encoding stage, respectively. Firstly, we concatenate features from the encoding and decoding stages along the channel dimension. Then, a convolution is used to reduce the channel counts and reduce the process’s computational costs to obtain two weights through our SKAF, which can be expressed as:
where represents Selective Kernel Attention Fusion Module, stands for the convolutional layer, and denotes the operation of concatenating features across the channel dimension. Next, we feed the two obtained weights into two branches for multiplication. Through this operator, we obtain the selected facial features from hybrid features obtained by fusing the encoding and decoding features. Finally, we add the features of the two branches:
@@
| Methods | PSNR/SSIM | VIF | LPIPS | Speed | Params |
| LGFI w/o SA | 27.75/0.7944 | 0.4652 | 0.1886 | 33ms | 12.3M |
| LGFI w/o RDFE | 27.51/0.7840 | 0.4495 | 0.2085 | 21ms | 7.2M |
| LGFI w/o SKAF | 27.74/0.7932 | 0.4611 | 0.1979 | 25ms | 8.21M |
| LGFI | 27.83/0.7961 | 0.4725 | 0.1821 | 34ms | 12.62M |
@@ III-D2 Perceptual loss
To enhance the visual quality of super-resolution images, we apply perceptual loss. This involves using a pre-trained VGG19 [39] model to extract facial features from both the HR images and our generated FSR images. Then, we compare the obtained perceptual features of HR and FSR images to constrain the generation of FSR features. Therefore, the perceptual loss can be described as:
@@ III-D3 Adversarial loss
GANs have been shown to be effective in reconstructing photorealistic images [37, 38]. GAN generates FSR results through the generator while using the discriminator to distinguish between ground truth and FSR results, which ultimately enables the generator to generate realistic FSR results in the process of constant confrontation. This process is denoted as:
@@ IV-A Datasets and Evaluation Metrics
We utilize the CelebA [4] dataset for training and evaluation on CelebA [4], Helen [40], and SCface [41] datasets, respectively. We center-crop the aligned faces and resize them to pixels to obtain HR images. These HR images are then downsampled to pixels using bicubic interpolation, producing the corresponding LR images. In our experiments, we randomly chose 18,000 CelebA images for training and 1,000 for testing. In addition, we also utilize the SCface test set as a real evaluation dataset. To measure the quality of FSR results, we use PSNR [42], SSIM [42], LPIPS [43], VIF [44], and FID [45].
@@ IV-B Implementation details
We implement our model using PyTorch on an NVIDIA GeForce RTX 3090. The network is optimized using the Adam optimizer, with parameters set to and . The initial learning rate is , with separate learning rates for the generator and discriminator set at and , respectively. The loss function weights are configured as , , and .
@@ IV-C1 Study of LGFI
LGFI is proposed to extract local features and global relationships of images, which represents a new attempt to interact with local and global information. To verify the reasonableness of our design of LGFI, as shown in Table I, we design four ablation models. The first model removes the SA, labeled “LGFI w/o SA”. The second model removes RDFE, labeled as “LGFI w/o RDFE”. The third model removes SKAF, labeled as “LGFI w/o SKAF”. We have the following observations: (a) Introducing SA and RDFE alone can improve model performance. This is because the above two modules can capture local and global features to promote facial feature reconstruction, including facial details and overall contours. (b) Model performance has been significantly increased by introducing the SKAF to capture the relationship between local and global facial features. This is because our SKAF can promote interaction between our SA and RDFE, integrating richer information and providing supplementary information for the final FSR.
Furthermore, we provide a visual comparison in Fig. 4, illustrating the impact of removing certain components from LGFI. The reconstructed images exhibit noticeable blurring or artifacts, highlighting the importance of these components. Additionally, Fig. 5 presents heatmap visualizations of the outputs from different components within LGFI, with their corresponding locations in the network indicated in Fig. 2. Specifically, SKAF effectively integrates the global contours captured by SA with the regional features extracted by RDEF. This integration enables the model to focus more on essential facial structures and components while reducing emphasis on less critical details, such as hair. In addition, we quantitatively evaluate the computational efficiency of each module by comparing inference latency and parameter counts in Table I, where the RDEF module has the greatest impact on both inference time and parameter counts. This is attributed to its multi-branch fusion strategy and deep refinement operations, which introduce additional computational complexity. However, RDEF delivers substantial performance improvements, which is overall worthwhile.
@@ IV-C2 Comparison between LGFI and Transformer
As shown in Fig. 6, LGFI uses a dual-branch structure to represent the local and global features. In contrast, the traditional Transformer in Restormer [50] uses a serial structure to link local and global features. To verify the effectiveness of LGFI, we replace all LGFIs with Transformers and conduct comparative experiments with similar parameters between the two models. From Table II, the network’s performance using LGFI is better when the two networks maintain similar parameters. This is because LGFI utilizes the features of both local and global branches for interaction, facilitating the communication of multiscale facial information.
@@ IV-C3 Comparison between RDFE and FFN
The feed-forward network (FFN) performs independent nonlinear transformations of the inputs at each position to help the Transformer capture local features, but it cannot extract multiscale features, which is not favorable for accurate FSR. In contrast, our RDFE can extract multiscale local features well. To compare RDFE and FFN, we replace RDFE with FFN while keeping the parameters of the two models similar. As shown in Table III, since FFN’s ability to capture feature interactions is limited compared to our RDFE, which utilizes multiple branches to capture different receptive field features, our RDFE performs better than FFN with similar computational cost.
@@ IV-C4 Effectiveness of RDFE
In RDFE, a three-branch network guided by an attention mechanism is used for deep feature extraction, and the feature refinement module is used to enrich feature representation. To verify the effectiveness of RDFE, we conduct multiple ablation experiments. We designed five improved models. The first model adopts a single branch structure of depthwise convolution, labeled as “Single path ( dw)”. The second model adopts a single branch structure of depthwise convolution, labeled as “Single path ( dw)”. The third model adopts a single branch structure of depthwise convolution, labeled as “Single path ( dw)”. The fourth model removes attention units labeled as ”w/o AU”. The fifth model removes the feature refinement module, labeled as ”w/o FRM”. From the Table IV, we have the following observations: (a) By comparing the first three rows and the last row of the table, it can be seen that multiscale branching facilitates the model’s performance due to its ability to extract face features at different levels; (b) From the comparison between the second and the last rows of the table and the last row, it can be seen that using attention units (AU) to guide three-branch feature extraction can enable the model to adaptively allocate weights, enhance the representation of important facial information, and thus improve model performance; (c) From the last two rows of the table, we can conclude that the feature refinement module (FRM) can further integrate multiscale information, refine multiscale fusion features, and thus improve performance.
@@ IV-C6 Study of EDFF
This section presents a set of experiments to validate the effectiveness of our EDFF, a module tailored for fusing multiscale features. We add EDFF to SPARNet [17], which uses EDFF to connect the encoding and decoding stages in SPARNet and send them to the next decoding stage. Additionally, we add EDFF to SFMNet [20], and the specific operation is the same as in SPARNet. From the results of Table VI, we can see that although the parameters of both models increase slightly, the performance of the models improves, which precisely proves that EDFF is helpful for feature fusion in the encoding and decoding stages.
| Methods | PSNR | SSIM | VIF | FID |
| FSRGAN [12] | 25.02 | 0.7279 | 0.3400 | 146.55 |
| DICGAN [14] | 25.59 | 0.7398 | 0.3925 | 144.25 |
| SPARGAN [17] | 25.86 | 0.7518 | 0.3932 | 149.54 |
| SFMGAN [20] | 25.96 | 0.7618 | 0.4019 | 141.23 |
| AMIGAN (Ours) | 26.35 | 0.7769 | 0.4101 | 122.43 |
@@ IV-D Comparisons with Other Methods
We compares AMINet with existing FSR methods, including SAN [46], RCAN [24], HAN [47], SwinIR [9], FSRNet [12], DICNet [14], FACN [48], SPARNet [17], SISN [18], AD-GNN [49], Restormer-M [50], LAAT [51], ELSFace [52], SFMNet [20] and SPADNet [53].
IV-D1 Comparisons on CelebA dataset
@@
IV-D2 Comparisons on Helen dataset
@@
IV-D3 Comparisons with GAN-based methods
We present AMIGAN as a novel approach to enhance the visual fidelity of image restoration. To demonstrate its effectiveness, we compare AMIGAN with existing GAN-based methods, including FSRGAN [12], DICGAN [14], SPARGAN [17], and SFMGAN [20]. In addition to conventional metrics, we adopt FID [45] for quantitative evaluation. Results on the Helen dataset (Table VIII) show that AMIGAN consistently outperforms prior methods. Visual comparisons in Fig. 10 further highlight AMIGAN’s superior ability to restore fine facial structures and texture details, particularly around the mouth and nose, delivering clearer and more realistic reconstructions with fewer artifacts.
IV-D4 Comparisons on Real-world surveillance faces
@@
All the above comparisons are tested on synthetic test sets, which fail to simulate real-world scenarios accurately. To further evaluate our model’s performance in real-world conditions, we also conduct experiments using low-quality face images from the SCface dataset [41]. As shown in Fig. 11, we compare the reconstruction results. From this figure, we find that the reconstruction results of face prior-based methods are not satisfactory. The challenge lies in accurately estimating priors from real-world LR facial images. Incorrect prior information can lead to misleading guidance during the reconstruction process. In contrast, our AMINet can restore clearer face details and faithful face structures. As shown in Table IX, we also provide a quantitative comparison of cosine similarity using the above methods. This result fully demonstrates our method’s effectiveness in real scenarios.
IV-E Model Complexity and Convergence Analysis
@@
In addition to the performance indicators mentioned earlier, the number of model parameters, inference time, and computational complexity are crucial factors in evaluating performance. As shown in the Table X, we have selected some models for comparison. Meanwhile, as shown in Fig. 1, we compare our model with existing ones in terms of parameters, PSNR values, and inference speed. We can see that AMINet can have faster inference time, smaller parameter count, and computational complexity with similar performance to SCTANet. In addition, we provide a small parameter version called ”AMINet-S”. AMINet-S performs similarly to SFMNet in terms of parameter quantity, while AMINet has more advantages in computational complexity and inference time. This is thanks to AMINet achieving high efficiency by promoting extensive multiscale feature exchange within the network. This design enables features from different receptive fields to interact effectively, allowing the model to select the most relevant information for facial reconstruction adaptively. Therefore, AMINet maintains strong FSR performance while balancing parameters and computational costs.
We also provide the training loss curve in Fig. 12, which shows a consistent decrease as iterations progress, indicating stable optimization and convergence. For clarity, the loss values are scaled by a factor of 100. We also visualize the intermediate results at different training stages, illustrating the clear improvement from early to late iterations as the model progressively learns more discriminative features. These observations collectively confirm the effectiveness of the training strategy and its stable convergence properties.
@@ V Discussion and Future Works
Future work will focus on enhancing AMINet’s robustness to extreme poses and challenging lighting conditions while accelerating inference. This includes integrating pose-invariant feature learning through attention-based mechanisms or 3D-aware priors for better side-profile restoration and developing adaptive illumination correction using physics-based relighting models or low-light enhancement. Additionally, optimizing the model with lightweight architectures, quantization, and efficient inference strategies will enable faster inference while maintaining high-quality restoration.